our work V:25.0 R-15 25.0
Answer #1

Solution) This problem is based on current electricity

Mainly two resistors in parallel

(1/Rp) = (1/R1) + ( 1/R2)

Rp = R1R2/(R1 + R2)

Two Resistors in series

Rs = R1 + R2

From ohms law we have V = iR

R1 and R2. in parallel, s) (25) Rey 2 3 ニ(1.02) (2s 3